OpenStudy (anonymous):

A long one :o. Myra wants to simplify the following expression. (2p − 7) + 2(7p + 9) Which option has the correct steps listed out for solving the expression?

OpenStudy (anonymous):

1) 2p − 7 + 2 + 7p + 2 + 9 using distributive property of addition 2) 2p + 7p – 7 + 2 + 2 + 9 simplifying 3) p (2 + 7) – 7 + 2 + 2 + 9 using commutative property of addition 4) 9p + 6 simplifying 1) 2p − 7 + 2 + 7p + 9 using distributive property of addition 2) 2p + 7p – 7 − 2 + 9 simplifying 3) p (2 + 7) + 0 using commutative property of addition 4) 9p simplifying 1) 2p − 7 + 2 (7p) + 9 using distributive property of addition 2) 2p + 14p −7 + 9 simplifying 3) p (2 + 14) −7 + 9 using commutative property of addition 4) 16p + 2 simplifying 1) 2p − 7 + 2 (7p) + 2 (9) using distributive property of addition 2) 2p + 14p – 7 + 18 using commutative property of addition 3) p (2 + 14) – 7 + 18 using distributive property of addition 4) 16p + 11 simplifying

OpenStudy (lalaly):

last one

OpenStudy (anonymous):

explain a lil bit please? :P

OpenStudy (lalaly):

lol well (2p-7)+2(7p+9) first bracket is not multiplied by anything so u can ignore it so 2p-7+2(7p+9) now using distributive property 2p-7+ 2(7p)+2(9) =2p-7+14p+18 =2p+14p+18-7 =16p+11

OpenStudy (anonymous):

Oh I see, TY :D!

OpenStudy (lalaly):

:)